OpenStudy (anonymous):

-5/8c = 20 A. 34 B. –32 C. –34 D. 32

OpenStudy (anonymous):

@i..am..legend

OpenStudy (unklerhaukus):

is that \[\frac{-5}{8c} = 20\]?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

get the c onto the other side, by multiplying both sides of the equation by c

OpenStudy (anonymous):

So -5 * 8? (-40)

OpenStudy (unklerhaukus):

wait a minute, is it real question \[\frac{-5}{8}c = 20\]

OpenStudy (anonymous):

Yes sorry

OpenStudy (unklerhaukus):

Well they are quite different equations. \[\frac{-5}{8}c = 20\] divide both sides of this equation by 5

OpenStudy (unklerhaukus):

can you do this @5EasyAnswers ?

OpenStudy (anonymous):

nooim stuck

OpenStudy (unklerhaukus):

\[\frac{-5}{8}c/5 = 20/5\\\frac{-\cancel5}{8}c\cancel{/5} = 20/5\]

OpenStudy (unklerhaukus):

-c/8 =20/5 can you divide the right hand side ?

OpenStudy (anonymous):

20/5 = 4

OpenStudy (unklerhaukus):

so you have -c/8 = 4 now multiply both sides of the equation by 8

OpenStudy (anonymous):

8 * 4 = 32

OpenStudy (unklerhaukus):

-c = 32 now what is c = ?

OpenStudy (anonymous):

32

OpenStudy (unklerhaukus):

minus c equal thirty-two so positive c equals ........ ......-...

OpenStudy (anonymous):

-32?

OpenStudy (unklerhaukus):

right

OpenStudy (anonymous):

Thank you